WRJPGCOM(1) WRJPGCOM(1) NAME wrjpgcom−inserttextcommentsintoaJPEGfile SYNOPSIS wrjpgcom[−replace][−commenttext][−cfilename][filename] DESCRIPTION wrjpgcomreadsthenamedJPEG/JFIFfile,orthestandardinputifnofileisnamed,andgeneratesanewJPEG/JFIFfileonstandardoutput.Acommentblockisaddedtothefile. TheJPEGstandardallows"comment"(COM)blockstooccurwithinaJPEGfile.Althoughthe standarddoesn’tactuallydefinewhatCOMblocksarefor,theyarewidelyusedtoholdusersuppliedtextstrings.Thisletsyouaddannotations,titles,indexterms,etctoyourJPEGfiles, andlaterretrievethemastext.COMblocksdonotinterferewiththeimagestoredintheJPEG file.ThemaximumsizeofaCOMblockis64K,butyoucanhaveasmanyofthemasyoulikein oneJPEGfile. wrjpgcomaddsaCOMblock,containingtextyouprovide,toaJPEGfile.Ordinarily,theCOM blockisaddedafteranyexistingCOMblocks;butyoucandeletetheoldCOMblocksifyouwish. OPTIONS Switchnamesmaybeabbreviated,andarenotcasesensitive. −replace DeleteanyexistingCOMblocksfromthefile. −commenttext SupplytextfornewCOMblockoncommandline. −cfilename ReadtextfornewCOMblockfromnamedfile. Ifyouhaveonlyonelineofcommenttexttoadd,youcanprovideitonthecommandlinewith −comment.Thecommenttextmustbesurroundedwithquotessothatitistreatedasasingle argument.Longercommentscanbereadfromatextfile. Ifyougiveneither−commentnor−cfile,thenwrjpgcomwillreadthecommenttextfrom standardinput.(InthiscaseaninputimagefilenameMUSTbesupplied,sothatthesource JPEGfilecomesfromsomewhereelse.)Youcanentermultiplelines,upto64KBworth.Type anend-of-fileindicator(usuallycontrol-D)toterminatethecommenttextentry. wrjpgcomwillnotaddaCOMblockiftheprovidedcommentstringisempty.Therefore −replace−comment""canbeusedtodeleteallCOMblocksfromafile. EXAMPLES Addashortcommenttoin.jpg,producingout.jpg: wrjpgcom−c"Viewofmybackyard"in.jpg>out.jpg Attachalongcommentpreviouslystoredincomment.txt: wrjpgcomin.jpg<comment.txt>out.jpg orequivalently wrjpgcom-cfilecomment.txt<in.jpg>out.jpg SEEALSO cjpeg(1),djpeg(1),rdjpgcom(1) AUTHOR IndependentJPEGGroup 30August1994 1